1. A MAN THROWS THE BRICKS TO A HEIGHT OF 12 METER WHERE THEY REACH WITH A SPEED OF 12 M/S. IF HE THROWS THE BRICKS SUCH THAT THEY JUST REACH THAT HEIGHT, WHAT PERCENTAGE OF ENERGY WILL BE SAVED?


a)29%


b)46%


c)38%


d)50%


2. A FORCE F=(2i + j + 3k)N ACTS ON A PARTICLE OF MASS 1 Kg FOR 2 seconds. IF INITIAL VELOCITY OF PARTICLE IS


u= (2i +j) m/s, speed of particle at the end of 2s will be


a)12m/s


b)6m/s


c)9m/s


d)4m/s


saumya shivhare

Grade:11

1 Answers

Chetan Mandayam Nayakar
13 years ago

if speesd at 12m is 12m/s and bricks travel a further distance before coming to rest 122=2gh,h=7.2, therefore,energy saved =7.2/19.2= 38%

v=u+at

v= (2+(2*2))i +(1+(2*1))j +(3*2)k, therefore speed2 =36+9+36=81,speed=9m/s